I need help please. Here is the problem: The measure of angle Q is (6x+16.8) and the measure of angle R is 2x. If angles Q and R are supplementary, what is the value of x? What are the measures for each angle?

Answer by nerdybill(7384)

The measure of angle Q is (6x+16.8) and the measure of angle R is 2x. If angles Q and R are supplementary, what is the value of x? What are the measures for each angle?
.
Since Q and R are supplementary:
R+Q = 180
2x + 6x+16.8 = 180
8x = 163.2
x = 20.4
.
Measure of R:
2x = 2(20.4) = 40.8 degrees
.
Measure of Q:
6x+16.8 = 6(20.4)+16.8 = 122.4+16.8 = 139.2 degrees